OpenStudy (anonymous):

What is the solution of log(2x - 3)125 = 3 ?

OpenStudy (luigi0210):

Rewrite it into an exponent

OpenStudy (anonymous):

what do you mean?

OpenStudy (luigi0210):

you know how log(3)x=2 is 3^2=x?

OpenStudy (anonymous):

125^(2x-3)=3? 5^3(2x-3)=3

OpenStudy (luigi0210):

(2x-3)^3=125

OpenStudy (anonymous):

so where would you go from here? sorry my book does not have any examples of this

OpenStudy (luigi0210):

take the cubic root of both sides

OpenStudy (anonymous):

2x-3=5?

OpenStudy (anonymous):

x=4

OpenStudy (luigi0210):

There you go :P

OpenStudy (anonymous):

can you please explain how you rearragned the problem?

OpenStudy (luigi0210):

\[\log_{x} k=y\] \[x^y=k\]

OpenStudy (anonymous):

thank you